# largesymbols"0D largesymbols"0D

### v3

@@ 2.2 The data
${\cal D}$ =$=\{\boldsymbol{x},\;\boldsymbol{y}\},$ (1)
[…]
@@
In some applications, the order of searches is also available, so $\boldsymbol{y}_{ij}$ can have a third element $y_{ij}^{(order)}$. However, because the search order is not always available, in this paper we choose not to make use of search order information (so $y_{ij}^{(order)}$ is not required). Conceptually, it is possible to pretrain a NNE that uses the search order information.
Our pretrained NNE assumes that $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$ and $\boldsymbol{x}_{i}^{(cons)}$ are all standardized to have zero means and unit variances. The means and variances here are computed across $i$ and $j$, e.g., the mean for $\boldsymbol{x}_{ij}^{(prod)}$ is $\frac{1}{nJ}\sum_{i=1}^{n}\sum_{j=1}^{J}\boldsymbol{x}_{ij}^{(prod)}$. If the attributes are not standardized, we standardize them before using the pretrained NNE to estimate $\boldsymbol{\theta}$. We then rescale the estimate of $\boldsymbol{\theta}$ to make it interpretable with respect to the original, non-standardized attributes.
@@
[…]
@@
A note about Table 1 is that no cap is imposed on $n$ for applying the pretrained NNE. But a cap on $n$ is necessary in pretraining, because we cannot simulate data of infinite size. This cap is currently set at $42e$ 4. Pretrained NNE is not directly applicable to data larger than this cap. However, if an application has larger data, we can split the data and average the estimates across the splits. [^2] The splitting and averaging are incorporated in our code at pnnehome.github.io.
Too few searches or purchases in data indicate difficulty for estimating $\boldsymbol{\theta}$. Therefore, we do not pretrain NNE to accommodate such data. Specifically, we look at:
@@
We require that the buy rate is at least 0.5% and the search rate is at least 1%. Again, these thresholds are configurable prior to pretraining. During pretraining, simulated datasets where search rates or buy rates are lower than thresholds are dropped.
As will be described later, the pretraining includes simulated datasets with non-normally distributed attributes. Nevertheless, the pretrained NNE may still perform poorly if certain attributes are too heavily skewed or have too many outliers. Therefore, we recommend using the usual practices to prepare data as in any empirical analysis. These include trimming, winsorizing, transforming variables, and dropping any attribute that causes (almost) collinearity.
@@ 3.1 Overview
Let ${\cal D}$ denote a generic dataset as defined in expression (1). Let $\boldsymbol{\theta}$ collect the parameters of the search model specified in Section 2.1. The task of NNE is to map ${\cal D}$ to an estimate of $\boldsymbol{\theta}$. It does so through two functions, $\boldsymbol{f}$ and $\boldsymbol{h}$.
@@
  • The first function $\boldsymbol{h}$ calculates a set of data patterns of ${\cal D}$, collected in vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$. This function is specified by us (rather than trained by examples).
@@
Conceptually, one could forgo specifying $\boldsymbol{h}$ and instead train a neural net mapping directly from ${\cal D}$ to $\boldsymbol{\theta}$. But the size of ${\cal D}$ is much larger than $\boldsymbol{m}$. In addition, the mapping from ${\cal D}$ to $\boldsymbol{\theta}$ tends to be much more complex than from $\boldsymbol{m}$ to $\boldsymbol{\theta}$. So the specification of function $\boldsymbol{h}$ substantially reduces the complexity of the learning task. The use of aggregate data patterns also allows privacy-preserving estimation, as discussed in Section 1.1.
In general, the vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$ may include any data pattern of ${\cal D}$. The foremost candidates are various data moments (this is why we use notation $\boldsymbol{m}$), such as means, variances, correlations. More sophisticated candidates are coefficients of reduced-form regressions. The idea of using reduced-form coefficients as “moments” originated from indirect inference (Gourieroux et al. 1993). Including reduced-form coefficients in $\boldsymbol{m}$ turns out to substantially improve the estimation accuracy of NNE, and is an important extension to Wei and Jiang (2024).
The detailed specification of $\boldsymbol{h}$ in our pretrained NNE is given in appendix. Below, we give a summary.
@@
  • To capture search decisions, we include the coefficients from logit regression that predicts $y_{ij}^{(search)}$. Similarly, to capture purchase decisions, we include the coefficients from multinomial logit regression that predicts $y_{ij}^{(buy)}$ conditional on search, i.e., $y_{ij}^{(search)}=1$. These regressions are at the product level.
  • We include the coefficients from a set of regressions similar to above, but at the consumer level (e.g., whether consumer $i$ bought anything, whether $i$ made any non-free search).
Next, we turn to the function $\boldsymbol{f}$, which is the core of the pretrained NNE. It is a neural net trained using simulated datasets. The training is conducted as follows.
  • We use the search model to simulate $L$ datasets ${\cal D}^{(1)}$, ${\cal D}^{(2)}$, …, ${\cal D}^{(L)}$, each under a different parameter, $\boldsymbol{\theta}^{(1)}$, $\boldsymbol{\theta}^{(2)}$, …, $\boldsymbol{\theta}^{(L)}$.
@@
We have left out three details in this training process: (i) the configuration of the neural net, (ii) the distribution from which we draw $\boldsymbol{\theta}^{(\ell)}$, and (iii) the distribution from which we draw the product, consumer, and advertising attributes in ${\cal D}^{(\ell)}$. These details are discussed in Section 3.2. At a high level, the distributions in (ii) and (iii) should be broad so that NNE can accommodate a variety of real datasets in applications, and yet not too broad so that NNE wastes resources learning scenarios that are unlikely to arise in reality. A good balance makes the training more effective for any given $L$. Meanwhile, it is important for us to point out the pretrained NNE’s estimates in applications (Section 4) are not sensitive to the specifications of details (i)-(iii).
@@
The description so far has focused on point estimates. For the standard error of $\widehat{\boldsymbol{\theta}}$, we recommend bootstrapping. This boils down to applying the pretrained NNE to resampled copies of ${\cal D}$. Because it incurs very little cost to apply pretrained NNE to each copy of data, bootstrapping standard errors is not computationally expensive at all. [^3] The bootstrapping here omits the uncertainty from the training of neural net. However, because of the large number of training examples ($L\simeq 1$ million), this uncertainty is small relative to the sampling uncertainty of ${\cal D}$.
A final issue in real applications is that some datasets may not be suitable for our pretrained NNE, because their true data generating processes are poorly described by the search model in Section 2.1. Actually, in such cases it is advisable not to apply this search model. For this reason, it will be useful to have a way to detect ill-suited datasets. We describe a detection method in the appendix. The basic idea is to check whether the real data patterns are an outlier compared to the data patterns used in pretraining.
@@ 3.2 Pretraining details
So far we have given an overview of how the pretrained NNE is constructed. Next, we provide more details on the training of the neural net $\boldsymbol{f}$. Readers more interested in the applications may skip these details without much loss of continuity.
@@ 3.2.1 The neural net
A neural net accepts fixed-length input and gives fixed-length output. But the search model parameter $\boldsymbol{\theta}$ has a dimension of $d+2$, where $d$ is the total number of attributes and it varies across different datasets. We address this issue through padding. Specifically, recall that $d\equiv d^{(prod)}+d^{(cons)}+d^{(ads)}$. Let subscript “max” denote the maximal value of a dimension. For example, Table 1 sets $d_{max}^{(prod)}=8$ and $d_{max}=8+5+2=15$. We configure the neural net to always output $d_{max}+2$ values. Out of these output values, only $d+2$ values are useful. The rest are the padded “idle” values. Using ellipses to represent idle values, we can write the output as:
$(\enspace\alpha_{0}\enspace,\enspace\boldsymbol{\alpha}^{\prime}\enspace,\underbrace{\enspace\cdot\cdot\cdot\enspace}_{d_{max}^{(ads)}-d^{(ads)}},\enspace\eta_{0}\enspace,\enspace\boldsymbol{\eta}^{\prime}\enspace,\underbrace{\enspace\cdot\cdot\cdot\enspace}_{d_{max}^{(cons)}-d^{(cons)}},\enspace\boldsymbol{\beta}^{\prime}\enspace,\underbrace{\enspace\cdot\cdot\cdot\enspace}_{d_{max}^{(prod)}-d^{(prod)}}).$
In training, the loss function ignores idle output values. In applications, the idle output values are disregarded.
A similar arrangement applies to the neural net’s input. The length of $\boldsymbol{m}$ varies by the numbers of attributes in the data. For our current settings in Table 1, the maximal length of $\boldsymbol{m}$ is 164. Any shorter $\boldsymbol{m}$ is padded with zeros to this maximal length before we feed it into the neural net. The exact detail is given in the specification of $\boldsymbol{m}$ in appendix.
We currently use $L\simeq$ 1e6 training examples (this is after dropping “corner” datasets with very low search or buy rates; see Section 2.2). We hold out $45e$ 4 examples for validation. Increasing $L$ further only marginally reduces the validation loss. We also use hold-out validation to choose hyperparameters of the neural net. One particular hyperparameter is the neural net architecture. We end up using three hidden layers with 512, 256, and 256 nodes. The pretrained NNE’s estimates in applications (Section 4) are not sensitive to the neural net architecture.

### v4

@@ 2.2 The data
${\cal D}$ $=\{\boldsymbol{x},\;\boldsymbol{y}\},$ (1)
$\text{where}:$ $\quad\boldsymbol{x}=\{\{\boldsymbol{x}_{ij}^{(prod)},\boldsymbol{x}_{ij}^{(ads)}\}_{j=1}^{J},\>\boldsymbol{x}_{i}^{(cons)}\}_{i=1}^{n},$
@@
In some applications, the order of searches is also observable, so $\boldsymbol{y}_{ij}$ can have a third element $y_{ij}^{(order)}$. However, because the search order is not always observable, in this paper we choose not to make use of search order information (so $y_{ij}^{(order)}$ is not required). Conceptually, it is straightforward to pretrain a separate NNE that uses the search order information.
Our pretrained NNE assumes that $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$ and $\boldsymbol{x}_{i}^{(cons)}$ are all standardized to have zero means and unit variances. The means and variances here are computed across $i$ and $j$, e.g., the mean for $\boldsymbol{x}_{ij}^{(prod)}$ is $\frac{1}{nJ}\sum_{i=1}^{n}\sum_{j=1}^{J}\boldsymbol{x}_{ij}^{(prod)}$. If the attributes are not standardized, we standardize them before using the pretrained NNE to estimate $\boldsymbol{\theta}$. We then rescale the estimate of $\boldsymbol{\theta}$ to make it consistent with respect to the original, non-standardized attributes.
@@
max 8 2 5 35  $\begin{array}[]{c}\infty\;\text{ (application)}\\ 2e4\text{ (pretraining)}\end{array}$
@@
A note about Table 1 is that no cap is imposed on $n$ for applying the pretrained NNE. But a cap on $n$ is necessary in pretraining, because we cannot simulate data of infinite size. This cap is currently set at $2e4$. The pretrained NNE is not directly applicable to data larger than this cap. However, in practice, we can split a very large dataset and apply the pretrained NNE to each split. We then take the average of the estimates across the splits.
It is difficult to estimate $\boldsymbol{\theta}$ if the data contain too few searches or purchases. Therefore, we do not pretrain NNE to accommodate such data. Specifically, we look at:
@@
We require that the buy rate is at least 0.5% and the search rate is at least 1%. Again, these thresholds are configurable prior to pretraining. During pretraining, simulated datasets where search rates or buy rates are lower than the thresholds are dropped.
As we will describe later, the pretraining incorporates non-normally distributed attributes. Nevertheless, the pretrained NNE may still perform poorly if certain attributes are too heavily skewed or have too many outliers. Therefore, we recommend using the usual practices to prepare data as in any empirical analysis. These include trimming, winsorizing, transforming variables, and dropping any attribute that causes (almost) collinearity.
@@ 3.1 Overview
Let $\cal D$} denote a generic dataset as defined in expression (1). Let $\boldsymbol{\theta}$ be the parameter vector of the search model specified in Section 2.1. The task of NNE is to map $\cal D$} to an estimate of $\boldsymbol{\theta}$. It does so through two functions, $\boldsymbol{f}$ and $\boldsymbol{h}$.
@@
  • The first function $\boldsymbol{h}$ calculates a set of data patterns of $\cal D$}, collected in vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$. This function is specified by us (rather than trained via examples).
@@
Conceptually, one could forgo specifying $\boldsymbol{h}$ and instead train a neural net mapping directly from $\cal D$} to $\boldsymbol{\theta}$. But the size of $\cal D$} is much larger than $\boldsymbol{m}$. In addition, the mapping from $\cal D$} to $\boldsymbol{\theta}$ tends to be much more complex than from $\boldsymbol{m}$ to $\boldsymbol{\theta}$. So the specification of function $\boldsymbol{h}$ substantially reduces the complexity of the learning task. Further, the use of aggregate data patterns also allows privacy-preserving estimation, as discussed in Section 1.1.
In general, the vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$ may include any data pattern of $\cal D$}. Ideally, we want to include data patterns that are most informative about $\boldsymbol{\theta}$ and yet inexpensive to compute. The foremost candidates are various data moments (hence the notation $\boldsymbol{m}$), such as means, variances, correlations. Further candidates are coefficients of reduced-form regressions. The idea of using reduced-form coefficients as “moments” originated from indirect inference (Gourieroux et al. 1993). Including reduced-form coefficients in $\boldsymbol{m}$ turns out to substantially improve the estimation accuracy of NNE, and is an important extension to Wei and Jiang (2024).
The detailed specification of $\boldsymbol{h}$ in our pretrained NNE is given in the appendix. Below, we give a summary.
@@
  • To capture search decisions, we include the coefficients from a logit regression that predicts $y_{ij}^{(search)}$. Similarly, to capture purchase decisions, we include the coefficients from a multinomial logit regression that predicts $y_{ij}^{(buy)}$ conditional on search, i.e., $y_{ij}^{(search)}=1$. These regressions are at the product level.
  • We include the coefficients from a set of regressions that predict outcomes at the consumer level (e.g., whether consumer $i$ bought anything, whether $i$ made any non-free search).
Next, we turn to the function $\boldsymbol{f}$, which plays a central role in the pretrained NNE. The function is a neural net trained using simulated datasets. The training is conducted as follows.
  • We use the search model to simulate $L$ datasets, ${\cal D}^{(1)}$, ${\cal D}^{(2)}$, …, ${\cal D}^{(L)}$, each under a different parameter vector, $\boldsymbol{\theta}^{(1)}$, $\boldsymbol{\theta}^{(2)}$, …, $\boldsymbol{\theta}^{(L)}$.
@@
We have left out three details in this training process: (i) the configuration of the neural net, (ii) the distribution from which we draw $\boldsymbol{\theta}^{(\ell)}$, and (iii) the distribution from which we draw the product, consumer, and advertising attributes in ${\cal D}^{(\ell)}$. These details are discussed in Section 3.3. Generally speaking, the distributions in (ii) and (iii) should be broad enough for NNE to accommodate a variety of real datasets in applications, but not so broad that NNE wastes resources learning scenarios that are unlikely in reality. A good balance makes the training more effective for a given $L$. Meanwhile, it is important for us to point out the pretrained NNE’s estimates in applications (Section 4) are not sensitive to the specifications of (i)-(iii).
@@
The description so far has focused on point estimates. For the standard error of $\widehat{\boldsymbol{\theta}}$, we recommend bootstrapping. This boils down to applying the pretrained NNE to resampled copies of $\cal D$}. Because applying the pretrained NNE to each copy of $\cal D$} incurs very little cost, bootstrapping standard errors is not computationally expensive at all. [^2] The bootstrapping here omits the uncertainty in $\boldsymbol{f}$ resulting from the neural net training. However, given the large number of training examples ($L\simeq 1$ million), this uncertainty tends to be small relative to the sampling uncertainty of $\cal D$}.
Finally, a practical issue is how NNE behaves when the search model is misspecified for the dataset in a real application. NNE takes data patterns as input and outputs $\widehat{\boldsymbol{\theta}}$. This output is given so that, when the search model is simulated at $\widehat{\boldsymbol{\theta}}$, the resulting data patterns are close to the input data patterns. Importantly, this behavior does not require the input data patterns to be generated by the search model. In this sense, NNE is well-behaved under misspecification. [^3] Also see Section 3.2 that shows the pretrained NNE converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$, which tends to be well-behaved under misspecification provided that the observed $\boldsymbol{m}$ is within the model-implied support of $\boldsymbol{m}$. However, this behavior does require the input data patterns to have positive probability under the search model and priors. If not, the input will lie outside the range of inputs seen in training and the neural net has to extrapolate. We call such cases “ill-suited” for the pretrained NNE, and it is useful to detect them in practice. We provide a detection method in the appendix.
3.2 Connection to Bayesian estimation
The eventual test for a pretrained model is to see how it performs in applications. We do so for the pretrained NNE in Section 4 and find that it provides accurate and sensible estimates. That said, from a theoretical standpoint, it is important to understand the econometric meaning of the pretrained NNE’s estimate. Below, we will explain that pretrained NNE has a familiar Bayesian interpretation, in that it calculates the posterior mean of $\boldsymbol{\theta}$ given $\boldsymbol{m}$. In words, it provides the best estimate of the structural model’s parameter $\boldsymbol{\theta}$ given the data patterns $\boldsymbol{m}$.
Continuing with the notation in (1), let ${\cal D}=\{\boldsymbol{x},\boldsymbol{y}\}$ denote a dataset. In particular, $\boldsymbol{x}$ and $\boldsymbol{y}$ collect the observations across all consumers. The vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$ denotes a set of data patterns. We start with a result with the standard (non-pretrained) NNE, studied in Wei and Jiang (2024). The main difference between standard NNE and pretrained NNE is that the former is trained for a specific real dataset. The training examples do not simulate $\boldsymbol{x}$ but directly use the observed $\boldsymbol{x}$ from that real dataset.
The standard NNE converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m},\boldsymbol{x})$ as $L\rightarrow\infty$ (see Wei and Jiang 2024). This result does not require large $n$ and holds for any value of $n$. Also note that we make it explicit that the expectation conditions on a specific $\boldsymbol{x}$. [^4] The notation of Wei and Jiang (2024) omits the conditioning on $\boldsymbol{x}$ because the paper assumes an underlying probability space that is already conditional on the observed $\boldsymbol{x}$. The $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m},\boldsymbol{x})$ is known as a limited-information Bayesian posterior mean (e.g., Kim 2002), where “limited-information” indicates that the posterior relies on less information than the full data. To compare, the full-information posterior is $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{y},\boldsymbol{x})$. Ideally, we would like to know the full-information posterior, but when computing it is difficult, the limited-information posterior provides a practical alternative.
We now turn to the pretrained NNE. In this case, the training examples do not assume a specific $\boldsymbol{x}$, but instead draw $\boldsymbol{x}$ from a distribution. Consequently, the estimate by the pretrained NNE converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$ as $L\rightarrow\infty$, where the expectation no longer conditions on $\boldsymbol{x}$.
We provide a formal statement for this convergence result. To be precise, we consider the joint probability distribution between $\boldsymbol{\theta}$ and $\boldsymbol{m}$, $P(\boldsymbol{\theta},\boldsymbol{m})$. This joint distribution is implied by the search model, our specification of $\boldsymbol{m}$, and the distributions from which we draw $\boldsymbol{\theta}$ and $\boldsymbol{x}$ in pretraining. Specifically, in terms of probability densities,
$p(\boldsymbol{\theta},\boldsymbol{m})=\int p(\boldsymbol{m}|\boldsymbol{\theta},\boldsymbol{x})p(\boldsymbol{\theta})p(\boldsymbol{x})\>d\boldsymbol{x}.$ (2)
Here, $p(\boldsymbol{m}|\boldsymbol{\theta},\boldsymbol{x})$ is the likelihood of observing data patterns $\boldsymbol{m}$ under the search model. Terms $p(\boldsymbol{\theta})$ and $p(\boldsymbol{x})$ are densities from which we draw $\boldsymbol{\theta}$ and $\boldsymbol{x}$ in pretraining, and $p(\boldsymbol{\theta},\boldsymbol{x})=p(\boldsymbol{\theta})p(\boldsymbol{x})$ as we will draw them independently. With the probability density in equation (2), we have
$\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})=\frac{\int\boldsymbol{\theta}p(\boldsymbol{\theta},\boldsymbol{m})d\boldsymbol{\theta}}{\int p(\boldsymbol{\theta},\boldsymbol{m})d\boldsymbol{\theta}}.$
Each training example $\{\boldsymbol{\theta}^{(\ell)},\boldsymbol{m}^{(\ell)}\}$ is a draw from $P(\boldsymbol{\theta},\boldsymbol{m})$. We use $\boldsymbol{f}_{L}$ to denote the neural net trained with $L$ training examples. That is, $\boldsymbol{f}_{L}=\min_{\boldsymbol{\zeta}\in{\cal F}_{L}}\sum_{\ell=1}^{L}\lVert\boldsymbol{\zeta}(\boldsymbol{m}^{(\ell)})-\boldsymbol{\theta}^{(\ell)}\rVert^{2}$, where ${\cal F}_{L}$ is the same class of single-hidden-layer neural nets defined in Wei and Jiang (2024). The result below states that the function $\boldsymbol{f}_{L}$ converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$ as a function of $\boldsymbol{m}$.
Proposition 1.
Suppose: (i) $\Theta$ is compact, (ii) $\boldsymbol{m}$ has a compact support, (iii) $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$ is a continuous function of $\boldsymbol{m}$. We have $\int\lVert\boldsymbol{f}_{L}(\boldsymbol{m})-\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})\rVert^{2}dP(\boldsymbol{m})\rightarrow 0$ in probability as $L\rightarrow\infty$.
The continuity condition should hold in most structural models as it only requires the posterior mean not to change abruptly for a slight change in the observed data patterns. In comparison, the compactness conditions can be more restrictive. While the compactness of $\Theta$ is common in econometric settings, it technically precludes unbounded priors such as the normal distribution. In practice, we can trim the prior at extreme values to ensure a compact $\Theta$. Similarly, the support-compactness of $\boldsymbol{m}$ can be satisfied if we trim extreme data patterns in training. Nevertheless, in applications (Section 4), we find that the pretrained NNE’s estimates show very little difference whether trimming is used in pretraining. Finally, it is worth emphasizing that the conditions in Proposition 1 are sufficient but not necessary.
Given this convergence result, a natural question is how to interpret $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$. Formally, it is another limited-information posterior mean. However, for a more intuitive perspective, consider a practical scenario where a researcher wants to apply the search model to a company’s data. For privacy reasons, the company provides the researcher with only aggregate data patterns $\boldsymbol{m}$, instead of the full data $\{\boldsymbol{x},\boldsymbol{y}\}$. In this case, the researcher’s best estimate for the search model parameter will be precisely $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$. From equation (2), we see that for $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$ to be well defined, the researcher should form a prior belief over $\boldsymbol{\theta}$, $p(\boldsymbol{\theta})$, and a prior belief over $\boldsymbol{x}$, $p(\boldsymbol{x})$. In the pretrained NNE, these priors are the distributions from which $\boldsymbol{\theta}$ and $\boldsymbol{x}$ are drawn during pretraining (details in Section 3.3).
@@ 3.3 Pretraining details
Section 3.1 gave an overview of our pretrained NNE, leaving out several details. We provide these details below. Readers interested mainly in the applications may skip these details without much loss of continuity.
@@ 3.3.1 The neural net
A neural net accepts fixed-length input and gives fixed-length output. But the search model parameter $\boldsymbol{\theta}$ has a size of $d+2$, where $d$ is the total number of attributes and it varies across different datasets. We address this issue through padding. Specifically, recall that $d\equiv d^{(prod)}+d^{(ads)}+d^{(cons)}$. Use subscript “max” to denote the maximum allowed value. For example, Table 1 sets $d_{max}^{(prod)}=8$ and $d_{max}=8+2+5=15$. We configure the neural net to always output $d_{max}+2$ values. Out of these output values, only $d+2$ values are useful. The rest are the padded “idle” values. Using ellipses to represent idle values, we can write the output as:
$(\kern 5.0pt\alpha_{0}\kern 5.0pt,\kern 5.0pt\boldsymbol{\alpha}^{\prime}\kern 5.0pt,\underbrace{\kern 5.0pt\cdot\cdot\cdot\kern 5.0pt}_{d_{max}^{(ads)}-d^{(ads)}},\kern 5.0pt\eta_{0}\kern 5.0pt,\kern 5.0pt\boldsymbol{\eta}^{\prime}\kern 5.0pt,\underbrace{\kern 5.0pt\cdot\cdot\cdot\kern 5.0pt}_{d_{max}^{(cons)}-d^{(cons)}},\kern 5.0pt\boldsymbol{\beta}^{\prime}\kern 5.0pt,\underbrace{\kern 5.0pt\cdot\cdot\cdot\kern 5.0pt}_{d_{max}^{(prod)}-d^{(prod)}}).$
During training, the idle output values are ignored by the loss function. In applications, the idle output values are discarded.
A similar arrangement applies to the neural net’s input. The length of $\boldsymbol{m}$ varies by the numbers of attributes in the data. For our current settings in Table 1, the maximal length of $\boldsymbol{m}$ is 164. Any shorter $\boldsymbol{m}$ is padded with zeros to this maximal length before we feed it into the neural net. The exact detail of $\boldsymbol{m}$ is given in the appendix.
We currently use $L\simeq$ 1e6 training examples (after dropping “corner” datasets with very low search or buy rates; see Section 2.2). We hold out $5e4$ examples for validation. Increasing $L$ further only marginally reduces the validation loss. We also use the hold-out validation to choose the neural net architecture. While Proposition 1 is formally stated with single-hidden-layer neural nets, in practice we find that using a few more hidden layers can offer lower validation losses. We end up using three hidden layers with 512, 256, and 256 nodes. The pretrained NNE’s estimates in applications (Section 4) are not sensitive to the neural net architecture.